OpenStudy (acacia):

Solve the equation of log↓a 16 = -4

OpenStudy (kirbykirby):

\(log_a 16=-4\) \(a^{-4}=16\) \[\frac{1}{a^4}=16\] \[a^4=\frac{1}{16}\] what do you think a could be?

OpenStudy (kirbykirby):

There's another way to solve it too

OpenStudy (anonymous):



OpenStudy (anonymous):

first way looks the best

OpenStudy (acacia):

I keep getting answers like 8 and 4

OpenStudy (kirbykirby):

\(log_x 16=-4\) \[\frac{log(16)}{log(x)}=-4\] \[log(x)=\frac{log(16)}{-4}\] \[x=10^{log(16)/-4}\]

OpenStudy (kirbykirby):

but this method is more cumbersome I find

OpenStudy (kirbykirby):

and requires a calculator

OpenStudy (kirbykirby):

\(a^4=16\)... is you put a=2 it works. But here, we have \[a^4=\frac{1}{16}\]

OpenStudy (acacia):

Ohhh I see now @kirbykirby

OpenStudy (kirbykirby):

:)